OpenStudy (anonymous):

305 degrees in radians in terms of pi

OpenStudy (anonymous):

pi = 180 degrees

OpenStudy (anonymous):

divide by 180, stick a \(\pi\) next to it

OpenStudy (unklerhaukus):

\[305°=305°\times 1=305°\times \frac{\pi[\text{radians}]}{180°}=\]